A 30.21 g sample of Ar (Argon) gas has a volume of 900 mL at a temperature of 200°C.
i. Calculate the pressure using the Ideal Gas Law.
ii. Calculate the corrected pressure using the Van der Waals Equation.
a=1.336 atm×L2 mol2 b=0.03201 L mol
1
Expert's answer
2020-04-18T07:07:37-0400

Molecular weight of Ar = 40 g/mole

30.21 g of Ar = 30.21/40 mole

=0.755 mole (n)

Volume,V = 900mL

=.9L

Temperature,T = 200°C

= 473K

Pressure,P = nRT/V where R=0.082

=32.54 atm


Now according to Vander Waals equation,

Pressure,P= (nRT/(V-nb))-(n2a/V2)

= 32.49 atm